# The Generic Temperature Response of Large Biochemical Networks

### v1

@@ Abstract
Biological systems are remarkably susceptible to relatively small temperature changes. The most obvious example is fever, when a modest rise in body temperature of only few Kelvin has strong effects on our immune system and how it fights pathogens. Another very important example is climate change, when even smaller temperature changes lead to dramatic shifts in ecosystems. Although it is generally accepted that the main effect of an increase in temperature is the acceleration of biochemical reactions according to the Arrhenius equation, it is not clear how it effects large biochemical networks with complicated architectures. For developmental systems like fly and frog, it has been shown that the system response to temperature deviates in a characteristic manner from the linear Arrhenius plot of single reactions, but a rigorous explanation has not been given yet. Here we use a graph theoretical interpretation of the mean first passage times of a biochemical master equation to give a statistical description. We find that in the limit of large system size and if the network has a bias towards a target state, then the Arrhenius plot is generically quadratic, in excellent agreement with experimental data for developmental times in fly and frog.
@@ I Introduction
High fever has a dramatic effect on our body, but from the physics point of view, it is only a modest change: increasing the human body temperature by three degrees is less than one percent on the absolute temperature scale [1]. An increase by the same amount due to global warming would most likely result in an extensive loss of biodiversity [2, 3]. As illustrated by these examples, complex biological systems are remarkably susceptible to changes in temperature. The explanation for these sensitive responses to temperature has been given already in the 19th century by Arrhenius, who suggested that the rates of all chemical reactions are exponentially increased by increased temperature [4]. This can be verified by an Arrhenius plot, in which the logarithm of the kinetic rate decreases in a linear fashion as a function of the inverse temperature $1/T$. This insight has been confirmed over and over again for single chemical reactions and has led to many important advances over the last century [5]. Moreover it has been made more rigorous by the theories by Kramers [6, 7] and Eyring [8, 9] on potential barriers as transition states in chemical reactions.
@@
There is one subject area for which the investigation of temperature effects has been relatively systematic and comprehensive, and that is the case of biochemical oscillators [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Here the general picture has emerged that they often come with compensation mechanisms which can assure that their functioning is unaltered within the physiological temperature range. A notable example for this general observation are circadian clocks, whose function is to instruct the organism about upcoming changes due to the diurnal cycle. Because temperature changes are one of the main consequences of changing solar input, circadian clocks are typically temperature-compensated, to ensure reliable time measurements [39, 40, 41, 42, 43]. Another important example of temperature compensation is chemotaxis of swimming organisms like E. Coli, which have to reliably find food sources despite temperature changes and gradients in their environment [47]. Although the existence of temperature compensation for biochemical oscillators and chemotaxis proves the relevance of temperature for biochemical networks, it does not instruct us about its generic effect, exactly because in these cases it is compensated by specific mechanisms, typically by the action of proteins that have evolved for that purpose.
@@
Because developmental systems are often shielded from temperature changes (e.g. due to an egg shell or a placenta which protects the embryo), they seem to have evolved less temperature compensation and therefore appear to be better model systems to study the effect of temperature changes on biochemical networks [48]. Measurements of the prepupal development time of Drosophila under different temperature conditions have been published almost one hundred years ago by Bliss [24]. The results clearly showed a concave curve in the Arrhenius plot, as opposed to the straight line usually found for single reactions. About a decade later, Powsner [25] obtained similar results for the embryonic, larval and pupal phase of the development of Drosophila, already hypothesizing that this may be a consequence of the complexity of the underlying network. More recently, Crapse et al. [26] published a study in which the embryonic development of Drosophila was subdivided into twelve phases. The authors measured the temperature dependence of the completion time of each of them. On this level, a concave temperature response emerged as well, which could be accurately described by a quadratic fit in the Arrhenius plot. In the same study, similar results were also found for Xenopus laevis embryos. The authors also performed simulations for a linear sequence of Poisson processes with statistically distributed parameters, but the results could not fully reproduce the shapes of the experimental curves [26]. Moreover, this approach disregards the fact that the underlying biochemical networks is bound to be highly non-linear, with complicated feedbacks and network motifs [49]. In general, although there has been some work on understanding the temperature dependence of development rates from first principles [50, 51, 52, 53, 54], the main body of work has remained empirically [55, 56, 57, 58, 59, 60].
Here we give a rigorous derivation why large biochemical networks typically show a quadratic shape in the Arrhenius plot. Our proof is based on concepts from stochastic dynamics and graph theory and is valid for all network architectures. Thus, our work solves the long-standing question what is the generic temperature response of large and complicated biochemical networks. We also show that our results are in excellent agreement with experimental data from developmental systems like fly and frog. Finally, our approach paves the way to deal with smaller biochemical systems and more specific temperature responses, including protein denaturation.
@@ II Results
II.1 Meaning of mean first passage time
A general framework to mathematically describe a chemical reaction network is to express it as time-dependent probability distribution $p_{i}(t)$ on $N+1$ discrete states, where without loss of generality we denote states $1$ and $N+1$ as start and target states, respectively. For a linear chain we thus would have $N$ reactions, each with forward and back directions, but in general, the system could contain many loops. The time evolution of $p_{i}$ is governed by the master equation[61]:
@@ II.1 Meaning of mean first passage time
Realistic biochemical networks typically exhibit a large number of intertwined reactions, meaning that $N\gg 1$ and Eq. (1) then describes a large system of coupled ordinary differential equations, for which it is difficult to find a full mathematical solution. Another important limitation is that often not all connections and/or rates might be known. Together this raises the question if one can make progress without explicitly solving the complete system.
Figure 1: The activation of cell division by an epidermal growth factor (EGF) via the mitogen-activated protein kinase (MAPK) signaling pathway is a biochemical network of central importance for developmental processes. Not all involved steps are shown in this cartoon (for more details see [62]).The mean first passage time $\langle\tau\rangle$ indicates how long it takes on average for an external growth stimulus to trigger a reaction in the nucleus. The guiding question in this work is to describe the temperature dependence of $\langle\tau\rangle$ in a many-component biochemical network under the assumption that the elementary reaction rates obey an Arrhenius-like temperature dependence.
Indeed, certain features of the system Eq. (1) can be computed without the need for a complete solution [63]. A notable example are first-passage times (also called first-hitting times or exit times) [64, 65] and in particular, the mean first passage time (MFPT), which characterizes the typical completion time of the process of interest. The main example in this work is a developmental process seen as the consequence of a complex network of biochemical reactions. Developmental systems are very large and complex and besides biochemistry also involve many spatial processes, such as cytoskeletal rearrangements [66]. However, experimental measurement of heat generation in zebrafish embryos combined with modelling of biochemical networks suggests that even in this case, the rate-limiting steps are in the biochemical control system for the cell cycle [67], which in itself is again a large and complex biochemical network [68]. To give an instructive example of a small and transparent biochemical network that is active during development and that can be described with the master equation from Eq. (1), in Fig. 1 we show the mitogen-activated protein kinase (MAPK) signaling pathway triggered by epidermal growth factor (EGF). In this case, the MFPT $\langle\tau\rangle$ describes the typical time for the signal to reach from the plasma membrane to the nucleus, where gene expression is changed. Here we ask who such MFPTs can be calculated for this and more complex networks, and how they depend on temperature.
@@ II.2.1 Some concepts from graph theory
It is often a useful approach to represent networks described by a master equation as weighted and directed graphs, i.e., vertices and directed edges (pairs of vertices), where the edges have a weight attributed to them [69]. Then the vertices are the states, the directed edges are the possible jumps and their weights are the jumping rates $k_{ij}$ from vertex $i$ to vertex $j$. In the following, graphs are always understood as weighted and directed. Also the rates $k_{ij}$ are simply referred to as edges, for the ease of notation, setting $k_{ij}=0$ if there is no transition from $i$ to $j$.
@@
(a) An example graph
(b) A spanning tree
(c) A spanning forest
Figure 2: The full directed graph on three vertices (a) with a spanning trees rooted at $3$ (b) and a spanning forest of two trees rooted at $2$ and $3$ (c).
@@
II.2.2 Graph theoretical interpretation of FPT moments
The standard text book approach to find first-passage times (FPTs) is to evaluate them in Laplace-space[70, 65, 64]. However, this becomes increasingly cumbersome for larger systems. Here we instead use a solution that can be derived with graph theory [71, 72]. Consider the master equation on a finite space $i=1,...N+1$ with time-independent rates $k_{ij}$. Then the formal solution to Eq. (2) is given by:
@@ II.2.2 Graph theoretical interpretation of FPT moments
By a generalization of Kirchhoff’s theorem to weighted and directed graphs[74], the determinant of $-K^{T}$ is given by:
@@
Figure 3: All spanning trees of the complete graph on four vertices rooted at $4$. The vertex labels are for all graphs as indicated for the first one.
$(-K^{T})^{-1}$ can be expessed as:
@@
where $M_{ji}$ is the determinant of the $(j,i)$-minor of $-K^{T}$, which can also be expressed as a sum with a graph theoretic interpretation[71]:
@@
where $\mathcal{F}_{[j,N+1]}^{i\rightarrow j}$ are the spanning forests with two trees of the graph, one rooted at $j$ and containing $i$ and the other one rooted at $N+1$.
Figure 4: The two-tree spanning forests $\mathcal{F}_{[1,4]}^{1\rightarrow 1}$, $\mathcal{F}_{[2,4]}^{1\rightarrow 2}$ and $\mathcal{F}_{[1,3]}^{1\rightarrow 3}$ of the complete graph on four vertices rooted at $4$. The vertex labels are for all graphs as indicated for the first one.
@@
Note that similar graph theoretical counting schemes also exist for the steady state distributions of ergodic networks[75, 76]. In appendix A we show how this formula can be used to derive MFPTs for two example networks for which the results are also known from Laplace transforms.
Overall, the graph theoretical approach results in three possible approaches to the FPTs of a homogeneous master equation. The first one consists in solving the matrix exponential as given in Eq. (9) or, alternatively, the corresponding differential equation:
$\dot{\vec{f}}=K\vec{f},$ (20)
with $\vec{f}_{0}$ as above. This is numerically feasible for systems of moderate size. If no full analytic solution is available, the moments of the FPT can be computed by either inverting $-K^{T}$ algebraically and using Eq. (12) or by solving the combinatorial problem in Eq. (18).
First-passage time distributions can be broad, especially when one considers the statistics of rare events, and higher moments are therefore often of interest[64]. For processes with a clear bias towards a final state, the distribution of the completion time tends to be more narrow and its mean characterizes it sufficiently well[77], which is likely to be the case for most developmental systems. Indeed, the development times of Drosophila embryos varies only by a few percent[78].
@@ II.3 Effect of temperature
Since most biological organisms have evolutionary adapted to a relatively fixed thermal environment, it is reasonable to rephrase the Arrhenius equation in terms of the rate constant at a reference temperature $T_{0}$, introducing $K_{i}=k_{i}(T=T_{0})$. $T_{0}$ describes the temperature at which the organism usually operates. For instance, for homeothermic animals, the body temperature would serve as the natural reference point, while for poikilothermic animals, it would be the typical temperature of the environment. $K_{i}$ can be seen as the standard rates, and a crucial assumption for the following analysis is that these standard rates are independent of the activation energies $E_{i}$.
@@
$k_{ij}=A_{ij}e^{-\frac{E_{ij}}{k_{B}T}}=k_{ij}(T=T_{0})e^{-\frac{E_{ij}}{k_{B}}(\frac{1}{T}-\frac{1}{T_{0}})}=K_{ij}e^{-\Delta\beta E_{ij}},$ (22)

### v2

@@ Abstract
Biological systems are remarkably susceptible to relatively small temperature changes. The most obvious example is fever, when a modest rise in body temperature of only few Kelvin has strong effects on our immune system and how it fights pathogens. Another very important example is climate change, when even smaller temperature changes lead to dramatic shifts in ecosystems. Although it is generally accepted that the main effect of an increase in temperature is the acceleration of biochemical reactions according to the Arrhenius equation, it is not clear how it affects large biochemical networks with complicated architectures. For developmental systems like fly and frog, it has been shown that the system response to temperature deviates in a characteristic manner from the linear Arrhenius plot of single reactions, but a rigorous explanation has not been given yet. Here we use a graph-theoretical interpretation of the mean first-passage times of a biochemical master equation to give a statistical description. We find that in the limit of large system size and if the network has a bias towards a target state, then the Arrhenius plot is generically quadratic, in excellent agreement with numerical simulations for large networks as well as with experimental data for developmental times in fly and frog. We also discuss under which conditions this generic response can be violated, for example for linear chains, which have only one spanning tree.
@@ I Introduction
High fever has a dramatic effect on our body, but from the physics point of view, it is only a modest change: increasing the human body temperature by three degrees is less than one percent on the absolute temperature scale [1]. An increase of the same amount due to global warming would most likely result in an extensive loss of biodiversity [2, 3]. As illustrated by these examples, complex biological systems are remarkably susceptible to changes in temperature. The explanation for these sensitive responses to temperature was already given in the 19th century by Arrhenius, who suggested that the rates of all chemical reactions exponentially increase with rising temperature [4]. This can be verified by an Arrhenius plot, in which the logarithm of the kinetic rate decreases linearly as a function of the inverse temperature $1/T$. This insight has been confirmed over and over again for single chemical reactions and has led to many important advances over the last century [5]. Moreover, it has been made more rigorous by the theories by Kramers [6, 7] and Eyring [8, 9] on potential barriers as transition states in chemical reactions.
@@
There is one subject area for which the investigation of temperature effects has been relatively systematic and comprehensive, and that is the case of biochemical oscillators [35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46]. Here, the general picture has emerged that they often come with compensation mechanisms which can assure that their functioning is unaltered within the physiological temperature range. A notable example for this general observation are circadian clocks, whose function is to instruct the organism about upcoming changes due to the diurnal cycle. Because temperature changes are one of the main consequences of changing solar input, circadian clocks are typically temperature-compensated, to ensure reliable time measurements [39, 40, 41, 42, 43]. Another important example of temperature compensation is chemotaxis of swimming organisms like E. Coli, which have to reliably find food sources despite temperature changes and gradients in their environment [21]. Although the existence of temperature compensation for biochemical oscillators and chemotaxis proves the relevance of temperature for biochemical networks, it does not instruct us about its generic effect, exactly because in these cases it is compensated by specific mechanisms, typically by the action of proteins that have evolved for that purpose.
Figure 1: The Arrhenius plot depends on network size. (a) The temperature dependence of the rate of a single biochemical reaction, such as ligand-receptor binding, usually follows the Arrhenius equation, resulting in a linear relationship in the Arrhenius plot (logarithmic rates against inverse temperature). (b) For networks of intermediate size, like the MAPK-signaling pathway shown here, the temperature response is more complex and depends on microscopic details. (c) In large biochemical networks, such as those governing the development from a Drosophila fertilized egg to a larva, a quadratic dependence emerges in the Arrhenius plot.
@@
Because developmental systems are often shielded from temperature changes (e.g. due to an egg shell or a placenta which protects the embryo), they seem to have evolved less temperature compensation and therefore appear to be better model systems to study the effect of temperature changes on large biochemical networks [47, 48]. Measurements of the prepupal development time of Drosophila under different temperature conditions were published almost one hundred years ago by Bliss [25]. The results clearly showed a concave curve in the Arrhenius plot, as opposed to the straight line usually found for single reactions. About a decade later, Powsner [26] obtained similar results for the embryonic, larval and pupal phases of the development of Drosophila, already hypothesizing that this may be a consequence of the complexity of the underlying network. More recently, Crapse et al. [24] published a study in which the embryonic development of Drosophila was subdivided into twelve phases. The authors measured the temperature dependence of the completion time of each of them. On this level, a concave temperature response emerged as well, which could be accurately described by a quadratic fit in the Arrhenius plot. In the same study, similar results were also found for Xenopus laevis embryos. The authors also performed simulations for a linear sequence of Poisson processes with statistically distributed parameters, but the results could not fully reproduce the shapes of the experimental curves [24]. However, this approach disregards the fact that the underlying biochemical networks are bound to be highly non-linear, with complicated feedback mechanisms and network motifs [49]. In general, although there has been some work on understanding the temperature dependence of development rates from first principles [50, 51, 52, 53, 54], the main body of work has remained empirical [55, 56, 57, 58, 59, 60].
The general question addressed in this work is illustrated in Fig. 1. The temperature dependence of a single reaction is usually described well by the Arrhenius equation, resulting in a linear relationship in the Arrhenius plot (Fig. 1a). However, when considering complex biological processes, e.g. the whole mitogen-activated protein kinase (MAPK) signaling pathway triggered by epidermal growth factors (EGF), that transmits extracellular signals to the nucleus and regulates differentiation [61], the linear relation is lost and microscopic details might govern the response (Fig. 1b). In the limit of very large systems, like whole organisms such as Drosophila developing from a fertilized egg to a larva [62], the overall rate usually approaches a quadratic dependence (Fig. 1c). Here we provide a rigorous mathematical derivation of why large biochemical networks typically show a quadratic shape in the Arrhenius plot. Our proof is based on calculating first passage times using a graph-theoretical decomposition of the network into spanning trees and forests. Our result is valid for all network architectures, including feedback loops, as long as several spanning trees stay relevant in the limit of large system size. Thus, our work solves the long-standing question of what the generic temperature response of large and complicated biochemical networks is. To support our analytical findings, we have developed a numerical procedure that generates networks of given sizes and with a bias towards an endpoint, and then uses the graph-theoretical interpretation of mean first-passage times to generate their distribution. This numerical procedure confirms the generic temperature response predicted by our theory. We also show that our results are in excellent agreement with experimental data from developmental systems like fly and frog. Finally we discuss the cases in which the generic response does not occur, including linear chains, which have only one spanning tree. Our approach paves the way to deal with smaller biochemical systems and more specific temperature responses, including protein denaturation.
@@ II Results
II.1 Meaning of mean first-passage time
A general framework to mathematically describe a chemical reaction network is to express it as a time-dependent probability distribution $p_{i}(t)$ on $N_{v}$ discrete states, where without loss of generality we denote states $1$ and $N_{v}$ as start and target states, respectively. For a linear chain, we would have $N=N_{v}-1$ reactions, each with forward and backward directions, while in general, the system could contain many loops. For our theoretical considerations, one needs both $N+1$ and $N$, thus we first use $N$ as main variable; for the numerical part coming later, it is more convenient to use $N_{v}=N+1$.
The time evolution of $p_{i}$ is governed by the master equation [63, 64]:
@@ II.1 Meaning of mean first-passage time
Realistic biochemical networks typically exhibit a large number of intertwined reactions, meaning that $N\gg 1$ and Eq. (1) then describes a large system of coupled ordinary differential equations, for which it is difficult to find a full mathematical solution. Another important limitation is that often not all connections and/or rates might be known. Together, this raises the question if one can make progress without explicitly solving the complete system.
Indeed, certain features of the system described by Eq. (1) can be computed without the need for a complete solution [65]. A notable example are first-passage times (also called first-hitting times or exit times) [66, 67] and in particular, the mean first-passage time (MFPT), which characterizes the typical completion time of the process of interest. The main example in this work is a developmental process seen as the consequence of a complex network of biochemical reactions. Developmental systems are very large and complex and besides biochemistry also involve many spatial processes, such as cytoskeletal rearrangements [68]. However, experimental measurements of heat generation in zebrafish embryos combined with modeling of biochemical networks suggest that even in this case, the rate-limiting steps are in the biochemical control system for the cell cycle [69], which in itself is again a large and complex biochemical network [70]. To give an instructive example of a transparent biochemical network that is active during development and that can be described with the master equation from Eq. (1), one can think of the mitogen-activated protein kinase (MAPK) signaling pathway triggered by epidermal growth factors (EGF) shown in Fig. 1b. In this case, the MFPT $\langle\tau\rangle$ describes the typical time for the signal to reach from the plasma membrane to the nucleus, where gene expression is changed. Here we ask how such MFPTs can be calculated for this and more complex networks, and how they depend on temperature.
@@ II.2.1 Some concepts from graph theory
It is often a useful approach to represent networks described by a master equation as weighted and directed graphs, i.e., vertices and directed edges (pairs of vertices), where the edges have a weight attributed to them [71]. Then the vertices are the states, the directed edges are the possible jumps and their weights are the jumping rates $k_{ij}$ from vertex $i$ to vertex $j$. In the following, graphs are always understood as weighted and directed. Also, the rates $k_{ij}$ are simply referred to as edges, for the ease of notation, setting $k_{ij}=0$ if there is no transition from $i$ to $j$.
@@
Figure 2: An illustration of the graph-theoretical decomposition of networks into spanning trees and forests used throughout this work. (a) The full directed graph on three vertices. (b) A spanning tree rooted at $3$. (c) A spanning forest of two trees rooted at $2$ and $3$.
@@
II.2.2 Graph-theoretical interpretation of FPT moments
The standard textbook approach to find first-passage times (FPTs) is to evaluate them in Laplace space [72, 67, 66]. However, this becomes increasingly cumbersome for larger systems. Here, we instead use a solution that can be derived with graph theory [73, 74, 75]. Consider the master equation on a finite space i=1,N+1i=1,...N+1 with time-independent rates $k_{ij}$. Then the formal solution to Eq. (2) is given by:
@@ II.2.2 Graph-theoretical interpretation of FPT moments
By a generalization of Kirchhoff’s theorem to weighted and directed graphs [77], the determinant of $K^{T}$ is given by:
@@
Figure 3: All spanning trees of the complete graph on four vertices rooted at $4$. The graph-theoretical interpretation of the sum in Eq. 14 consists in finding the spanning trees. The vertex labels are for all graphs as indicated for the first one.
$(-K^{T})^{-1}$ can be expressed as:
@@
where $M_{ji}$ is the determinant of the $(j,i)$-minor of $K^{T}$, which can also be expressed as a sum with a graph-theoretical interpretation [73]:
@@
where $\mathcal{F}_{[j,N+1]}^{i\rightarrow j}$ are the spanning forests with two trees of the graph, one rooted at $j$ and containing $i$ and the other one rooted at $N+1$. Fig. (4) shows $\mathcal{F}_{[i,4]}^{1\rightarrow i}$ for the complete graph on 4 vertices.
Figure 4: All spanning forests of the complete graph on four vertices rooted at $4$. The graph-theoretical interpretation of the summands in Eq.16 consists in finding the two-tree spanning forests $\mathcal{F}_{[1,4]}^{1\rightarrow 1}$, $\mathcal{F}_{[2,4]}^{1\rightarrow 2}$ and $\mathcal{F}_{[3,4]}^{1\rightarrow 3}$ of the complete graph on four vertices rooted at $4$. The vertex labels are for all graphs as indicated for the first one.
@@
Note that similar graph-theoretical counting schemes also exist for the steady state distributions of ergodic networks [78, 79]. In Appendix A, we show how this formula can be used to derive MFPTs for two example networks for which the results are also known from Laplace transforms.
Overall, the graph-theoretical approach results in three possible approaches to the FPTs of a homogeneous master equation. The first one consists in solving the matrix exponential as given in Eq. (9) or, alternatively, the corresponding differential equation:
$\dot{\vec{f}}=K^{T}\vec{f},$ (20)
with $\vec{f}_{0}$ as above. This is numerically feasible for systems of moderate size. If no full analytic solution is available, the moments of the FPT can be computed by either inverting $K^{T}$ algebraically and using Eq. (12) or by solving the combinatorial problem in Eq. (18).
First-passage time distributions can be broad, especially when considering the statistics of rare events, and higher moments are therefore often of interest [66]. For processes with a clear bias towards a final state, the distribution of the completion time tends to be more narrow and its mean characterizes it sufficiently well [80], which is likely to be the case for most developmental systems. Indeed, the development times of Drosophila embryos vary only by a few percent [81].
@@ II.3 Effect of temperature
Since most biological organisms have evolutionarily adapted to a relatively fixed thermal environment, it is reasonable to rephrase the Arrhenius equation in terms of the rate constant at a reference temperature $T_{0}$, introducing $k^{0}_{i}=k_{i}(T=T_{0})$. $T_{0}$ describes the temperature at which the organism usually operates. For instance, for endothermic animals, the body temperature would serve as the natural reference point, while for exothermic animals, it would be the typical temperature of the environment. The $k^{0}_{i}$ can be seen as the standard rates, and a crucial assumption for the following analysis is that these standard rates are independent of the activation energies $E_{i}$.
@@
$k_{ij}$ $=A_{ij}e^{-\frac{E_{ij}}{k_{B}T}}=\underbrace{k_{ij}(T=T_{0})}_{=:k^{0}_{ij}}e^{-\frac{E_{ij}}{k_{B}}(\frac{1}{T}-\frac{1}{T_{0}})},$ (22)